# Torsional Carroll Gravity

### v1

@@ Abstract
The ultra-relativistic (Carrollian) regime of gravity has recently emerged as a fertile framework for exploring holography, non-Lorentzian symmetries, and geometric limit of General Relativity. In this letter, we establish the presence of a non-vanishing torsion within three-dimensional Carrollian gravity by constructing the Carrollian Mielke–Baekler (C-MB) gravity theory in its Chern–Simons formulation, obtained as the ultra-relativistic limit of the relativistic Mielke–Baekler model. The resulting C-MB theory features non-zero temporal torsion and curvature, together with spatial curvature, providing the most general three-dimensional Carrollian gravity model with these properties. Several known ultra-relativistic gravity theories arise as particular limits of this framework, highlighting its unifying character.
@@ I Introduction
At the level of gravity, the Carrollian theory arises from sending the speed of light to zero in the Einstein–Hilbert action [21]. In three spacetime dimensions, it can alternatively be formulated as a Chern–Simons (CS) theory based on the Carroll algebra [22, 23, 24, 25], the ultra-relativistic contraction of the Poincaré algebra [26]. The three-dimensional CS framework offers a powerful arena for probing structural aspects of gravity, including black-hole physics and thermodynamics [27]. Matter couplings to Carroll gravity have also been successfully constructed [28, 29, 30] performing systematic Carrollian expansions of relativistic models.
By contrast, the role of torsion in the ultra-relativistic regime remains poorly understood. While torsion is well established in Newton–Cartan gravity, arising naturally from gauging the Schrödinger algebra [31] and appearing in contexts such as Lifshitz holography [32] and the Quantum Hall Effect [33], no fully general formulation of torsional Carroll gravity exists. Current Carrollian models either enforce vanishing torsion or impose curvature constraints, preventing a complete geometric description. In particular, no three-dimensional Carrollian gravity theory simultaneously accommodating non-trivial torsion and curvature has been constructed to date, leaving torsional effects in the ultra-relativistic regime essentially unexplored.
In this letter, we fill this gap by constructing the first three-dimensional Carrollian gravity theory that simultaneously exhibits non-vanishing torsion and curvature. Starting from the relativistic Mielke-Baekler (MB) model [34, 35], a topological gravity theory that naturally incorporates torsion, we perform its ultra-relativistic contraction to obtain a Carrollian Mielke-Baekler (C-MB) theory in the CS formulation. The resulting C-MB model features non-zero temporal Carroll torsion, temporal curvature, and spatial curvature, thereby providing the most general and structurally complete Carrollian gravity theory to date. We further show that diverse known ultra-relativistic gravity models arise as specific limits of our construction, revealing its unifying character. This work places torsional Carroll gravity on solid geometric foundations and opens a new arena for exploring torsional effects in non-Lorentzian geometry and ultra-relativistic gravitational dynamics.

### v2

@@ Abstract
The ultra-relativistic (Carrollian) regime of gravity has recently emerged as a fertile framework for exploring holography, non-Lorentzian symmetries, and geometric limit of General Relativity. In this letter, we establish the presence of a non-vanishing torsion within three-dimensional Carrollian gravity by constructing the Carrollian Mielke-Baekler (C-MB) gravity theory in its Chern-Simons formulation, obtained as the ultra-relativistic limit of the relativistic Mielke-Baekler model. The resulting C-MB theory features non-zero temporal torsion and curvature, together with spatial curvature, providing the most general three-dimensional Carrollian gravity model with these properties. Temporal torsion affects non-affinity of null generators and boundary dynamics. Several known ultra-relativistic gravity theories arise as particular limits of this framework, highlighting its unifying character.
@@ I Introduction
At the level of gravity, the Carrollian theory arises from sending the speed of light to zero in the Einstein-Hilbert action [13]. In three spacetime dimensions, it can alternatively be formulated as a Chern-Simons (CS) theory based on the Carroll algebra [42, 48, 1, 23], the ultra-relativistic contraction of the Poincaré algebra [41]. The three-dimensional CS framework offers a powerful arena for probing structural aspects of gravity, including black-hole physics and thermodynamics [9]. Matter couplings to Carroll gravity have also been successfully constructed [35, 8, 10] performing systematic Carrollian expansions of relativistic models.
By contrast, the role of torsion in the ultra-relativistic regime remains poorly understood. While torsion is well established in Newton-Cartan gravity, arising naturally from gauging the Schrödinger algebra [11] and appearing in contexts such as Lifshitz holography [20] and the Quantum Hall Effect [19], no fully general formulation of torsional Carroll gravity exists. Current Carrollian models either enforce vanishing torsion or impose curvature constraints, preventing a complete geometric description. In particular, no three-dimensional Carrollian gravity theory simultaneously accommodating non-trivial torsion and curvature has been constructed to date, leaving torsional effects in the ultra-relativistic regime essentially unexplored.
In this letter, we fill this gap by constructing the first three-dimensional Carrollian gravity theory that simultaneously exhibits non-vanishing torsion and curvature. Starting from the relativistic Mielke-Baekler (MB) model [43, 3], a topological gravity theory that naturally incorporates torsion, we perform its ultra-relativistic contraction to obtain a Carrollian Mielke-Baekler (C-MB) theory in the CS formulation. The resulting C-MB model features non-zero temporal Carroll torsion, temporal curvature, and spatial curvature, thereby providing the most general and structurally complete Carrollian gravity theory to date. We further show that diverse known ultra-relativistic gravity models arise as specific limits of our construction, revealing its unifying character, and discuss some relevant physical implications of our C-MB theory on null hypersurfaces and boundary dynamics. This work places torsional Carroll gravity on solid geometric foundations and opens a new arena for exploring torsional effects in non-Lorentzian geometry and ultra-relativistic gravitational dynamics.
@@ III Carrollian Mielke-Baekler Gravity
III.1 Physical implications of torsional Carroll gravity
Carrollian gravity theories are known to play a central role in the description of null hypersurfaces and of boundary geometries in UR and asymptotically flat spacetimes. The C-MB model constructed in this Letter presents an intrinsic temporal torsion. We now highlight the physical implications of this geometric structure, focusing on its impact on the geometry of Carrollian horizons and on the structure of boundary dynamics, relevant for, e.g., flat holography. In particular, we show that temporal Carrollian torsion acquires a direct physical interpretation in terms of horizon non-affinity and defines a new geometric datum at null infinity.
III.1.1 Effects of Carrollian torsion on null hypersurfaces
Carrollian geometry naturally emerges as the intrinsic geometry of null hypersurfaces, such as black hole event horizons or null infinity [27, 39, 21, 17]. In this framework, the Carrollian clock 1-form $\tau$ is associated with the null generators of the hypersurface, while the spatial vielbein $e^{a}$ encodes the geometry of its spatial sections.
In the presence of temporal Carrollian torsion, this structure naturally endows null hypersurfaces with an additional geometric charge contribution (in our C-MB theory, the torsional boundary datum/source $q$), fixed by the underlying Carrollian symmetry data.
The temporal Carrollian torsion $T_{\mathrm{T}}:=d\tau+\epsilon^{ab}e_{a}\omega_{b}$ measures the failure of the null generators to be affinely parametrized. This is the Carrollian analogue of the standard relation
$k^{\nu}\nabla_{\nu}k^{\mu}=\kappa k^{\mu}$ (30)
for null generators $k^{\mu}$, with $T_{\text{T}}$ here playing the role of a geometric realization of the non-affinity $\kappa$. The torsion deformations, and their UR effects, are encoded in the covariant derivative acting on null generators. In the C-MB theory, the field equations impose
$T_{\mathrm{T}}=-\frac{q}{2}\epsilon^{ab}e_{a}e_{b},$ (31)
so that the non-affinity of the null generators is fixed by the torsional coupling $q$.
This provides a geometric and symmetry-based Carrollian interpretation of the surface gravity of non-extremal horizons, independent of the existence of a specific global black hole solution. In this sense, the parameter $q$ controls torsional contributions associated with horizon acceleration, while the parameter $p$ governs the intrinsic curvature of the spatial horizon sections. The C-MB theory therefore endows null hypersurfaces with a torsional charge, analogous, at the level of boundary data, to a chemical potential for horizon acceleration.
III.1.2 Implications on boundary dynamics
Carrollian geometries provide the natural boundary data for asymptotically flat spacetimes, where the dual dynamics is governed by Carrollian field theories with BMS-type symmetries. Within the C-MB framework, intrinsic temporal Carrollian torsion enlarges the space of admissible boundary geometries beyond the torsion-free case, while remaining compatible with a CS variational principle.
From the CS formulation, the on-shell variation of the action reduces to a boundary contribution,
$\delta I_{\text{C-MB}}\big|_{\text{on-shell}}=\int_{\partial\mathcal{M}}\langle\delta A\wedge A\rangle,$ (32)
where $\mathcal{M}$ is the three-dimensional bulk manifold and $\partial\mathcal{M}$ its boundary. Consistent boundary conditions and the associated boundary symplectic structure are therefore entirely determined by the invariant bilinear form (20).
From the latter, the boundary symplectic structure induced by the CS action identifies the canonical pairs $(\omega,\tau)$ and $(\omega^{a},e_{a})$, with symplectic weights proportional to $\tilde{\sigma}_{1}$. In particular, the non-vanishing pairing $\langle JH\rangle=-\tilde{\sigma}_{1}$ implies that the Carrollian clock 1-form $\tau$ is canonically conjugate to spatial rotations at the boundary, while the pairing $\langle K_{a}P_{b}\rangle=\tilde{\sigma}_{1}\delta_{ab}$ governs the boost-translation sector.
The variation of the surface charge associated with a Lie-algebra-valued gauge parameter $\lambda$ generating infinitesimal gauge transformations of the CS connection takes the standard CS form
$\delta Q[\lambda]=\int_{\partial\mathcal{M}}\langle\lambda\,\delta A\rangle.$ (33)
In the C-MB theory, the temporal Carrollian torsion constraint (31) enters the boundary variational principle as a fixed background source. As a consequence, the surface charge associated with Carrollian time translations acquires an explicit $q$-dependent contribution, persisting even in the absence of bulk curvature.
For generic values of $q$, the residual gauge transformations preserving the torsional boundary condition therefore form a deformation of the standard Carrollian (or $\mathfrak{bms}_{3}$-type) algebra. This deformation persists in the flat limit $p=0$, indicating that intrinsic temporal Carrollian torsion defines a genuinely new sector of Carrollian boundary dynamics. The parameter $p$, which controls spatial curvature, further allows for curved Carrollian boundary geometries, providing a unified framework encompassing flat, AdS-Carroll, and torsional Carrollian holography.
@@ IV Discussion
The C-MB theory modifies the topological description of null hypersurfaces in the UR regime, by extending Carrollian horizon models to include both curvature and intrinsic temporal torsion. Here, torsion affects non-affinity of null generators and acts as a geometric source for horizon acceleration. This framework naturally enables the study of torsion-dependent modifications of horizon symmetries, conserved charges, and near-horizon dynamics. As an illustration, one may consider the UR limit of the MB BTZ solution. Detailed derivations of Carrollian black holes with torsion, boundary conditions, surface charges, and the resulting deformed Carrollian symmetry algebra are left for future work.